A spherical tank 4.5m in diameter contains water to a depth of 3m. Find the volume of the water. 


1
Expert's answer
2021-12-17T10:48:06-0500

Volume of a spherical cap = 1/3 πh²(3R-h), where h=height of the spherical cap, 

      R= radius of the sphere.

In this question, R= 4.5÷2=2.25m, 

               h= 4.5-3=1.5m  

               π=22/7

Volume of water = volume of sphere - volume of cap

      = 4/3πr³ - ⅓πh²(3R-h)

      =4/3 π × 2.25³ -⅓π×1.5²(3×2.25-1.5)

      = 47.7129 - 12.3700

      = 35.3429